A firm has a Cobb Douglas production function given as

q=AL

α K β

a.

Solve for the factor demand functions

b.

If the firms’ competitive output price is p find the wage rate

c.

What is the share of the firms revenue paid to labour and

capital?

d.

If α =0.6, β =0.2 and A=1 find the LR labour and capital

demand curve equations


1
Expert's answer
2021-10-03T14:14:28-0400

"q=AL^\u03b1K^\u03b2"

a. Let the price of the output «q» be p.

Total cost of the firm "TC=wL+rK"

Profit "\u03a0 = pq -wL -rK"

"\u03a0 = pAL^\u03b1K^\u03b2 -wL -rK"

Profit is maximized when "\frac{∂Π}{∂L} = 0 \; and \; \frac{∂Π}{∂K} =0"

"αpAL^{α-1}K^β -w = 0 \;\;\;(1) \\

βpAL^αK^{β-1} -r =0 \;\;\;(2)"

Dividing two equations, we get:

"\frac{α}{β} \frac{K}{L} = \frac{w}{r} \\

wL = \frac{α}{β}Kr"

Substituting this in C = wL +rK, we get:

"\frac{α}{β}rK +rK=C \\

rK[\frac{α+β}{β}] =C \\

K^* = [\frac{β}{α +β} \times \frac{C}{r}] \\

L^* = [\frac{α}{α +β} \times \frac{C}{w}]"

b. From (1), we get:

"\u03b1pAL^{\u03b1-1}K^\u03b2 =w"

So, the wage rate w is

"w=\u03b1pAL^{\u03b1-1}K^\u03b2"

c. Total wage paid to labour "= wL = \u03b1pAL^{\u03b1}K^\u03b2"

Total sum paid to capital "= rK = \u03b2pAL^\u03b1K^{\u03b2}"

Share of firms revenue paid to labour "= \frac{wL}{pq}"

"= \frac{αpAL^{α}K^β}{pAL^{α}K^β} = α"

Share of firms revenue paid to capital "= \frac{rK}{pq}"

"= \frac{βpAL^αK^{β}}{pAL^{α}K^β}=β"

d. α=0.6, β=0.2, A=1

"K^* = \frac{0.2}{0.2+0.6} \times \frac{C}{r} = 0.25 \frac{C}{r} \\

L^* = \frac{0.6}{0.2+0.6} \times \frac{C}{w} = 0.75 \frac{C}{w}"